# ClustRecNet: A Novel End-to-End Deep Learning Framework for Clustering Algorithm Recommendation

### v2

@@ 2.3 Deep Model Architecture
2.4 Implementation Details
The described network architecture was chosen after an extensive experimentation with multiple network variants as it provided the best overall performance across our synthetic data repository. Moreover, the ablation study conducted on real-world data (see Section 3.2) showed that all components of the proposed network architecture are essential for generalizing across diverse clustering scenarios. Unlike traditional image-based CNNs, our model operates directly on tabular data encoded as 2D arrays, where rows correspond to individual data points (objects) and columns to features. We treat each dataset as a single-channel ”image” of shape $(1,N,D)$, where $N$ is the number of objects (padded to a fixed height), and $D$ is the number of features. This tabular treatment allows the convolutional network to learn local feature patterns across both object and feature dimensions.
@@ 2.4 Implementation Details
Figure 3: Normalized confusion matrices for 10 clustering algorithms generated by the proposed model on synthetic test data.
To capture long-range interactions, we incorporated a self-attention block into the model architecture, which follows the residual layers. Query and key projections reduce the channel dimension using 1×1 convolutions (with a reduction factor of 8), while the value projection maintains the input channels. Attention weights are computed across flattened spatial positions and applied back to the input with a residual connection. The output is flattened and passed through two fully connected layers - the first compresses the representation to 16 neurons, while the second produces predictions over clustering algorithms. The final model was trained using the Adam optimizer (learning rate $10^{-4}$, weight decay $7\times 10^{-3}$) with the Binary Cross-Entropy (BCE) loss with logits. Training proceeds by 30 epochs with a batch size of 32, using 10-fold cross-validation. All experiments were conducted on an NVIDIA A100-SXM4-40GB GPU. All implementation details, including source code and configuration, are available at our GitHub.
@@
3 Results and Discussion
3.1 Experiments with Synthetic Data
In this section, we rigorously evaluate ClustRecNet that uses as input an unlabeled $objects\times features$ dataset and outputs the most suitable algorithm(s) for its clustering, among those described in Section 2.2. During training, we used 90% of the data for 10-fold cross-validation to ensure robustness and reduce overfitting, with the remaining 10% reserved for final model validation. Additionally, weight decay regularization was applied through the Adam optimizer to further prevent model overparameterization and enhance generalization performance.
The BCE loss consistently decreased, indicating effective learning, from an initial point of 0.528 to 0.360 for training, and from 0.457 to 0.370 for validation. This reduction translates into F1-score improvements: For training, the F1-score increased from 0.534 to 0.763, while for validation it increased from 0.635 to 0.752, reflecting the model’s growing prediction accuracy. The Hamming distance decreased from 0.295 to 0.177 for training, and from 0.249 to 0.186 for validation, marking, respectively, a 40% and 25.3% reduction in misclassification.
To evaluate the effectiveness of our recommendation framework, we compared its performance on the test data against the following popular CVIs: Silhouette, CH, DB, and Dunn. We performed a grid search over predefined thresholds for each CVI to determine the one that yields the highest F1-score when compared to binary labels derived from ARI. The best selected threshold was then used to choose the most suitable clustering algorithm. The obtained F1-score, Hamming distance, and ARI results are reported in Table 1.
To compute the ARI, the clustering algorithm, selected either by our model or by a CVI, was executed using the hyperparameter configurations obtained via grid search, as detailed in Supplementary Material. Our model provided a mean F1-score of 0.757 and a mean ARI of 0.878, outperforming the leading CVIs (Silhouette and CH) by more than 5% in terms of F1-score and by 0.497 in terms of ARI. The mean Hamming distance was reduced by 53% compared to Silhouette, confirming the superiority of our method in predicting suitable clustering algorithms. We also carried out the Wilcoxon signed-rank test to assess whether the improvements achieved by our proposed model over traditional CVIs were statistically significant. The test results confirm the superiority of our model, with the obtained $p$-values less than 0.005 across all three evaluation metrics, strongly rejecting the null hypothesis of equal performance.
To provide a more detailed breakdown of prediction outcomes across 10 clustering algorithms considered in our study, we present their confusion matrices (see Figure 3). All reported values are normalized, making them directly comparable across algorithms. As shown, our proposed model consistently identifies the most suitable algorithms with high precision and recall, particularly using Spectral Clustering, Agglomerative Clustering, DBSCAN, HDBSCAN, and OPTICS, each exhibiting an accuracy over 80%.
@@ 3.2 Experiments with Real Data
Table 2: Results of our ablation study on real-world data presented in terms of ARI.
To assess the effectiveness of the proposed recommendation framework in practical settings, we conducted experiments on 10 well-known real-world datasets from the UCI Machine Learning Repository [^1] https://archive.ics.uci.edu. These datasets, summarized in Table 3, exhibit a diverse range of characteristics, including the number of objects, features, and classes, encompassing a variety of domains and complexities. Since all these datasets include ground-truth labels, we were able to compute ARI for each clustering solution obtained to compare the performance of our recommendation model to those of state-of-the-art AutoML-based clustering methods, including ML2DAC, AutoCluster, and AutoML4Clust (AML4C).
For each real-world dataset considered, our model was first applied to predict the most suitable clustering algorithm. The selected algorithm was then executed with default scikit-learn parameters over a predefined range of candidate cluster counts. Either Silhouette or CH index was then used to determine the optimal number of clusters. Such an approach simulates a typical unsupervised clustering scenario, where neither ground-truth labels nor the number of clusters are available at inference time.
To isolate the contribution of each architectural component of our model, we performed an ablation study in which one of its main components - either the CNN block, or two ResNet blocks, or the attention mechanism (ATN) - was removed, one at a time. The results, summarized in Table 2, highlight the importance of all three main components working in synergy. The full CH-based model provided the highest mean ARI of 0.2295 and a median ARI of 0.1775. In contrast, the average ARI across all ablated models dropped to 0.1850 (for mean) and to 0.1021 (for median), corresponding to losses of 21.5% and 73.8%, respectively. Compared to the closest CH-based ablation, the improvement in mean was rather modest, i.e. 2.8%, but the median gain was 44.4%, suggesting a much more stable clustering performance. Similarly, the full Silhouette-based model outperformed the best ablation by 26.1% in terms of the mean ARI and by 62.6% in terms of the median ARI, confirming the cumulative value of all model’s components.
To conduct the clustering performance comparison on real-world data, we first considered a baseline CNN using the first CNN block of our model architecture. This baseline shared all other elements of the pipeline – including training data, cluster estimation, and evaluation setup, allowing for a fair performance comparison. We further benchmarked our model against state-of-the-art AutoML-based clustering recommendation systems. Specifically, we carried out the official, containerized implementation of ML2DAC (available on GitHub [^2] https://github.com/tschechlovdev/ml2dac) – endorsed with a reproducibility badge by SIGMOD – on the same real-world data. Furthermore, we evaluated two built-in meta-learning approaches from the same container: AutoCluster and AML4C, covering a range of meta-learning strategies. A comprehensive summary of all comparative results is presented in Table 3.
Dataset Objects Features Classes ClustRecNet Baseline CNN AutoCluster AML4C ML2DAC
@@
Breast tissue 106 9 6 0.2682 0.2682 0.2682 0.2398 0.2398 0.2398 0.2099 0.2099
Ecoli 336 7 8 0.5011 0.5011 0.5011 0.5011 0.5011 0.5011 0.1535 0.3863
Glass 214 9 7 0.2453 0.2453 0.2453 0.1399 0.1399 0.1399 0.0994 0.1472
Haberman 306 3 2 0.1040 0.1040 0.0274 0.1556 0.0621 0.0621 0.0323 0.0323
Iris 150 4 3 0.5681 0.5681 0.5681 0.5584 0.5438 0.5438 0.5681 0.5681
Parkinsons 195 22 2 0.0514 0.1218 0.0514 0.0000 0.0000 0.0000 0.0473 0.0473
Transfusion 748 4 2 0.0193 0.0193 0.0131 0.0174 0.0304 0.0095 0.0196 0.0344
Vehicle 846 18 4 0.1097 0.0869 0.0878 0.0125 0.0119 0.0088 0.1380 0.1333
Vertebral column 310 6 3 0.3669 0.3669 0.1361 0.3669 0.0967 0.3369 0.1211 0.2885
Wine quality red 1599 11 6 0.0611 -0.0029 0.0157 0.0000 0.0000 0.0000 0.0505 0.0505
@@
Table 3: Detailed comparison of clustering algorithm recommendation methods across 10 real-world datasets in terms of ARI.
@@
Table 4: Wilcoxon signed-rank test $p$-values reflecting the performance improvement provided the proposed ClustRecNet model (CH variant) over the basic CNN and AutoML approaches on real-world data.
Moreover, we analyzed the ARI score distribution using a boxplot diagram (see Figure 4) and carried out the Wilcoxon signed-rank test to assess the performance improvement provided by ClustRecNet over the other competing methods (see Table 4). We found that the new model demonstrated statistically significant ARI improvements, with the Wilcoxon test $p$-values << 0.05, over the state-of-the-art AutoML-based clustering frameworks, including AutoCluster (CH, Silhouette, and DB variants), AML4C, and ML2DAC, as well as over the baseline CNN. Overall, our model systematically yielded the best results in terms of the mean and median ARI, regardless the cluster validity index (CH or Silhouette) used to determine the number of clusters, highlighting the model’s stability across different CVIs.
While the proposed model demonstrated strong performance and stability, several limitations remain. First, although the model generalizes well to real-world datasets, its predictive accuracy depends on the diversity and coverage of synthetic data used for training. Expanding the training set to better reflect edge-case scenarios, such as highly imbalanced clusters or very noisy data, could further improve its robustness. Second, our current pipeline selects the number of clusters using internal validation indices (CH or Silhouette in our experiments), while all of them are known to be imperfect. Incorporating into it a learned cluster count estimator or an ensemble of estimators may offer more reliable performance in the future. Third, the model assumes a tabular input representation; extending it to accommodate graph-based or time-series clustering problems could broaden its applicability.
Figure 4: Boxplot comparison of clustering algorithm recommendation methods across 10 real-world datasets.
@@ 4 Conclusion
We developed ClustRecNet – a novel recommendation framework to identify the most suitable clustering algorithms for a given dataset, showing that deep learning models can be effectively used to tackle this challenging issue. First, we generated a variety of synthetic datasets, employing different data generation strategies. Second, we applied 10 popular clustering algorithms on each of the datasets to evaluate their performance. The suitability of clustering algorithms for each dataset was assessed using ARI and encoded as a binary vector, convenient for multi-label classification tasks. Third, we designed and tested a DL-based hybrid network architecture with a CNN block, two residual blocks, and an attention mechanism. Such an architecture allows us to extract both low-level and high-level data features, with the attention mechanism highlighting critical features.
@@
Future research could build upon this framework in several directions. First, integrating into the model pipeline dimensionality reduction techniques, such as principal component analysis or autoencoders, could further optimize the system’s performance in high-dimensional spaces, particularly in scenarios where the feature set is large or sparsely populated. Additionally, exploring alternative feature extraction methods, such as graph neural networks, may enhance the model’s ability to capture complex relationships within data, especially for datasets with inherent hierarchical or graph-based structure. Another promising avenue is extending our framework to optimize the parameter settings for each algorithm recommended, possibly using reinforcement learning, thus widening its applicability and enhancing precision.
Supplementary Material
Hyperparameters Used with Clustering Algorithms
To ensure a fair and comprehensive comparison across clustering algorithms, we systematically explored a curated set of hyperparameter configurations for each method, while preserving computational tractability. The selected configurations aim to balance representational diversity and practical feasibility in automated clustering scenarios.
For the $k$-means and $k$-medians algorithms, we used the default k-means++ initialization strategy, increasing the number of random initializations (n_initn\_init) to 100 to enhance stability, while leaving other hyperparameters unchanged. For hierarchical algorithms, including Ward’s method and AC, we employed the nearest neighbors graph as the connectivityconnectivity parameter, varying the number of neighbors within the range $[3,10]$ with an increment of 1. In Ward’s method, the linkagelinkage criterion was fixed to ward, while in AC the three following alternative criteria were explored: complete, average, and single. To ensure a comprehensive range of distance measures, the following values of the metricmetric parameter were considered in AC: euclidean, l1, l2, manhattan, and cosine.
Regarding density-based methods, DBSCAN was executed using an epseps value ranging from 0.1 to 1.5 with an increment of 0.1. In HDBSCAN, the min_samplesmin\_samples parameter was selected from the range [3, 10] with a step of 1, while the min_cluster_sizemin\_cluster\_size parameter varied from 5 to 25 with a step of 5. Similarly, in OPTICS, we varied the min_samplesmin\_samples parameter as in HDBSCAN, min_cluster_sizemin\_cluster\_size varied from 0.1 to 0.5 with a step of 0.05, and $xi$ varied from 0.05 to 0.25 with a step of 0.05.
SC was evaluated by selecting the eigen_solvereigen\_solver parameter among arpack and lobpcg, and setting the affinityaffinity parameter either to nearest neighbors or rbf. In the BIRCH algorithm, the thresholdthreshold for subcluster identification ranged from 0.1 to 1.5 with an increment of 0.1. In GMM, the following four covariance types were examined: full, tied, diagonal, and spherical, allowing for flexible modeling of cluster shape and orientation.
All remaining algorithm parameters were left at their scikit-learn default values. The applied strategy ensures a reasonable balance between optimization quality and computational complexity for each clustering algorithm considered.
Train and Validation Convergence
Figure 5 in this Supplementary Material depicts the average Binary Cross-Entropy (BCE) loss (a), F1-score (b), and Hamming distance (c) results across all folds, obtained using the proposed ClustRecNet model. The trends in training and validation loss, F1-score, and Hamming distance provide insights into the model’s convergence behavior and generalization capacity. The BCE loss consistently declines as the number of epochs grows, indicating stable optimization. Training loss dropped from 0.528 to 0.360, while validation loss decreased from 0.457 to 0.370. This learning progression was reflected in a marked improvement in F1-scores: from 0.534 to 0.763 on training data, and from 0.635 to 0.752 on validation data, thus demonstrating increasing alignment with ground truth clustering labels. Simultaneously, the Hamming distance — a metric quantifying binary misclassification — showed an important reduction: from 0.295 to 0.177 for training and from 0.249 to 0.186 for validation. They correspond to relative improvements of approximately 40% and 25%, respectively. A close alignment between the training and validation curves across all metrics indicates strong generalization and stable convergence behavior.

### v3

@@ 2.5 Implementation Details
Unlike traditional image-based CNNs, our model operates directly on tabular data encoded as 2D arrays, where rows correspond to individual data points (i.e. objects) and columns to features. We treat each dataset as a single-channel ”image” of shape $(1,N,D)$, where $N$ is the number of objects (padded to a fixed height), and $D$ is the number of features. This tabular treatment allows the convolutional network to learn local feature patterns across both object and feature dimensions.
@@
To capture long-range interactions, we incorporated into the model architecture a self-attention block, which follows the residual layers. Query and key projections reduce the channel dimension using 1×1 convolutions (with a reduction factor of 8), while the value projection maintains the input channels. Attention weights are computed across flattened spatial positions and applied back to the input with a residual connection. The output is flattened and passed through two fully connected layers - the first compresses the representation to 16 neurons, while the second produces predictions over clustering algorithms. The final model was trained using the Adam optimizer (learning rate $10^{-4}$, weight decay $7\times 10^{-3}$). The recommendation task is formulated as a multi-label classification problem using the Binary Cross-Entropy (BCE) loss with logits. This formulation, which accommodates the inherent structural ambiguity of unsupervised learning, identifies a curated set of high-performing clustering algorithms without imposing a strict single-winner constraint. The training proceeds by 30 epochs with a batch size of 32, using 10-fold cross-validation. All experiments were conducted on an NVIDIA A100-SXM4-40GB GPU. All implementation details, including source code and configuration, are available on our GitHub.
2.6 Train and Validation Convergence
Figure 3 depicts the average Binary Cross-Entropy (BCE) loss (a), F1-score (b), and Hamming distance (c) results across all folds, obtained using the proposed ClustRecNet model.
The trends observed in the training and validation loss, F1-score, and Hamming distance results provide insights into the model’s convergence behavior and generalization capacity. The BCE loss consistently declines as the number of epochs grows, indicating stable optimization. The training loss drops from 0.528 to 0.360, while validation loss decreases from 0.457 to 0.370. These patterns are reflected in a marked improvement of the F1-score: from 0.534 to 0.763 on training data, and from 0.635 to 0.752 on validation data, thus demonstrating increasing alignment with ground truth clustering labels. At the same time, the Hamming distance — a metric used to quantify binary misclassification — shows a progressive decrease: from 0.295 to 0.177 for training, and from 0.249 to 0.186 for validation. These results correspond to relative improvements of approximately 40% and 25%, respectively. A close alignment between the training and validation curves across all metrics highlights strong generalization capacity and stable convergence behavior of ClustRecNet.
3 Results and Discussion
3.1 Experiments with Synthetic Data
In this section, we rigorously evaluate ClustRecNet that uses as input an unlabeled dataset, encoded as an $object\times feature$ matrix, and outputs the most suitable algorithm(s) for its clustering (i.e. for clustering its objects), among those described in Section 2.2. During training, we used 90% of the data for 10-fold cross-validation to ensure robustness and to reduce overfitting, with the remaining 10% reserved for model validation. Additionally, weight decay regularization was applied through the Adam optimizer to further prevent model overparameterization and to enhance generalization performance.
The BCE loss consistently decreased, indicating effective learning, from an initial point of 0.528 to 0.360 for training, and from 0.457 to 0.370 for validation. This reduction translates into the following F1-score improvements: for training, the F1-score increased from 0.534 to 0.763, and for validation it increased from 0.635 to 0.752, reflecting the model’s growing prediction accuracy. The Hamming distance decreased from 0.295 to 0.177 for training, and from 0.249 to 0.186 for validation, marking, respectively, a 40% and 25.3% reduction in misclassification.
To evaluate the effectiveness of our recommendation framework, we compared its performance on the test data against the following popular CVIs: Silhouette, CH, DB, and Dunn. We performed a grid search over predefined thresholds for each CVI to determine which one yields the highest F1-score when compared to binary labels derived from ARI. The best selected threshold was then used to choose the most suitable clustering algorithm. The obtained F1-score, Hamming distance, and ARI results are reported in Table 2.
@@ 3.1 Experiments with Synthetic Data
To compute the ARI value, the clustering algorithm, selected either by our model or by a CVI, was executed using the hyperparameter configurations obtained via grid search, as detailed in Section 2.3. Our model provided a mean F1-score of 0.757 and a mean ARI of 0.878, outperforming the leading CVIs (Silhouette and CH) by more than 5% in terms of F1-score and by 0.497 in terms of ARI. The mean Hamming distance was reduced by 53% compared to Silhouette, confirming the superiority of our model in predicting suitable clustering algorithms.
Moreover, we carried out the Wilcoxon signed-rank test to assess whether the improvements over traditional CVIs, achieved by ClustRecNet, were statistically significant. The test results confirm the superiority of our model, with the obtained $p$-values ¡ 0.005 across all three evaluation metrics, strongly rejecting the null hypothesis of equal performance.
To provide a more detailed breakdown of prediction outcomes across 10 clustering algorithms considered in our study, we present their confusion matrices (see Figure 4). All reported values are normalized, making them directly comparable across clustering algorithms. As shown, our proposed model consistently identifies suitable clustering algorithms with high precision and recall, particularly using Spectral Clustering, Agglomerative Clustering, DBSCAN, HDBSCAN, and OPTICS, each exhibiting an accuracy over 80%.
Figure 4: Normalized confusion matrices for the 10 clustering algorithms considered in our study, generated by ClustRecNet on synthetic test data.
@@ 3.2 Experiments with Real Data
To assess the effectiveness of the proposed recommendation framework in practical settings, we conducted experiments on 10 well-known real-world datasets from the UCI Machine Learning Repository [^2] https://archive.ics.uci.edu. These datasets, summarized in Table 3, exhibit a diverse range of dimensional characteristics and encompass a variety of domains and complexities. They include popular Ecoli, Glass, Iris, and Wine quality datasets, used in numerous unsupervised and supervised classification studies.
Table 3: Summary of real datasets from the UCI repository, including the number of objects, features, and classes in each dataset.
Dataset Objects Features Classes
Breast tissue 106 9 6
Ecoli 336 7 8
Glass 214 9 7
Haberman 306 3 2
Iris 150 4 3
Parkinsons 195 22 2
Transfusion 748 4 2
Vehicle 846 18 4
Vertebral column 310 6 3
Wine quality red 1599 11 6
Since the ground-truth labels are available for all these datasets, we were able to compute ARI for each clustering solution obtained to compare the performance of our recommendation model to those of state-of-the-art AutoML-based clustering methods, including ML2DAC, AutoCluster, and AutoML4Clust (AML4C).
For each real-world dataset reported in Table 3, our model was first applied to predict the most suitable clustering algorithm. The selected algorithm was then executed with default scikit-learn parameters over a predefined range of candidate numbers of clusters. Either the Silhouette or CH index was then used to determine the optimal number of clusters. Such an approach simulates a typical unsupervised clustering scenario, where neither ground-truth labels nor the number of clusters are known at inference time.
To isolate the contribution of each architectural component of our model, we performed an ablation study in which one of its main components - either the CNN block, or two ResNet blocks, or the attention mechanism (ATN) - was removed, one at a time. The results, summarized in Table 4, highlight the importance of all three main components working in synergy. The full CH-based model provided the highest mean ARI of 0.2295 and a median ARI of 0.1775. In contrast, the average ARI across all ablated models dropped to 0.1850 (for mean) and to 0.1021 (for median), corresponding to losses of 21.5% and 73.8%, respectively. Compared to the closest CH-based ablation, the improvement in mean was rather modest, i.e. 2.8%, but the median gain was 44.4%, suggesting a much more stable clustering performance. Similarly, the full Silhouette-based model outperformed the best ablation by 26.1% in terms of the mean ARI and by 62.6% in terms of the median ARI, confirming the cumulative value of all components of the model.
Table 4: Results of our ablation study presented in terms of ARI, conducted on the real-world UCI benchmarks from Table 3.
@@
To conduct a clustering performance comparison on real-world data, we first considered a baseline CNN using the first CNN block of our model architecture. Such a baseline shares all other elements of the pipeline – including training data, cluster estimation, and evaluation setup, allowing for a fair performance comparison. We further benchmarked our model against state-of-the-art AutoML-based clustering recommendation systems. Specifically, we used the official, containerized implementation of ML2DAC (available on GitHub [^3] https://github.com/tschechlovdev/ml2dac) – endorsed with a reproducibility badge by SIGMOD – on the same real-world data. Furthermore, we evaluated two built-in meta-learning approaches from the same container: AutoCluster and AML4C, covering a range of meta-learning strategies. The detailed comparative results are reported in Table 5.
Table 5: Detailed comparison of clustering algorithm recommendation methods, in terms of ARI, across the real-world UCI benchmarks from Table 3. The best per-dataset and the best average performances are highlighted in bold.
Dataset ClustRecNet Baseline CNN AutoCluster AML4C ML2DAC
@@
Breast tissue 0.2682 0.2682 0.2682 0.2398 0.2398 0.2398 0.2099 0.2099
Ecoli 0.5011 0.5011 0.5011 0.5011 0.5011 0.5011 0.1535 0.3863
Glass 0.2453 0.2453 0.2453 0.1399 0.1399 0.1399 0.0994 0.1472
Haberman 0.1040 0.1040 0.0274 0.1556 0.0621 0.0621 0.0323 0.0323
Iris 0.5681 0.5681 0.5681 0.5584 0.5438 0.5438 0.5681 0.5681
Parkinsons 0.0514 0.1218 0.0514 0.0000 0.0000 0.0000 0.0473 0.0473
Transfusion 0.0193 0.0193 0.0131 0.0174 0.0304 0.0095 0.0196 0.0344
Vehicle 0.1097 0.0869 0.0878 0.0125 0.0119 0.0088 0.1380 0.1333
Vertebral column 0.3669 0.3669 0.1361 0.3669 0.0967 0.3369 0.1211 0.2885
Wine quality red 0.0611 -0.0029 0.0157 0.0000 0.0000 0.0000 0.0505 0.0505
@@
Moreover, we analyzed the ARI distributions using a boxplot diagram (see Figure 5) and carried out the Wilcoxon signed-rank test to assess the performance improvement provided by ClustRecNet over the other competing methods (see Table 6). We found that the ClustRecNet demonstrated statistically significant ARI improvements, with the Wilcoxon test $p$-values << 0.05, over state-of-the-art AutoML-based clustering frameworks, including AutoCluster (CH, Silhouette, and DB variants), AML4C, and ML2DAC, as well as over the baseline CNN. Overall, our model systematically yielded the best results in terms of the mean and median ARI, regardless the cluster validity index (CH or Silhouette) used to determine the number of clusters, highlighting ClustRecNet’s stability across different CVIs.
Table 6: Wilcoxon signed-rank test $p$-values reflecting the performance improvement provided by the proposed ClustRecNet model (CH variant) over the basic CNN and state-of-the-art AutoML approaches; the test was conducted on the real-world UCI benchmarks from Table 3.
@@
Figure 5: Boxplots comparing clustering algorithm recommendation methods across the real-world UCI benchmarks from Table 3.
To demonstrate the robustness of our model, we conducted a comprehensive permutation invariance analysis across 10 real-world benchmark datasets, evaluating the impact of shuffling both row (objects) and column (features) orders of the input matrix. The aggregated results reveal that the model maintains high structural stability; notably, for half of the tested datasets, the model exhibited perfect invariance. Furthermore, the performance metrics remained highly consistent across permutations. Specifically, the mean ARI for row permutations was $0.2128\pm 0.0194$, while column permutations yielded a mean ARI of $0.2174\pm 0.0185$. Compared to the ClustRecNet mean ARI of $0.2295$ (calculated on the original input order), these minor fluctuations represent a negligible relative deviation ($6\%$). These empirical findings confirm that through intensive row-and-column shuffling during the training phase, the model achieves empirical permutation equivariance.
While the proposed ClustRecNet model demonstrated strong performance and stability, it also has a few limitations. First, although the model generalizes well to real-world datasets, its predictive accuracy depends on the diversity and coverage of synthetic data used for training. Expanding the training set to better reflect edge-case scenarios, such as highly imbalanced clusters or very noisy data including outliers, could further improve its robustness. Second, our current pipeline selects the number of clusters using an internal cluster validation index (CH or Silhouette in our experiments), while all of them are known to be imperfect [27]. Incorporating a learned cluster count estimator or an ensemble of estimators may offer more reliable performance in the future. Third, the model assumes a tabular input representation; extending it to accommodate graph-based or time-series input data could broaden its applicability.
@@ 4 Conclusion
We developed ClustRecNet – a novel recommendation framework to identify the most suitable clustering algorithms for a given tabular dataset, proving that deep learning can be effectively used to tackle this challenging issue. First, we generated a diverse synthetic data repository, employing well-established data generation strategies used in clustering. Second, we carried out 10 popular clustering algorithms on each dataset from this repository, assessing the quality of the obtained clustering solutions. The suitability of clustering algorithms for each dataset was evaluated using ARI, and encoded as a binary vector convenient for multi-label classification tasks. Third, we designed and tested a DL-based hybrid network architecture with a CNN block, two residual blocks, and an attention mechanism. Such an architecture allowed us to extract both low-level and high-level data features, with the attention mechanism highlighting critical features.
@@
Future research could build upon this framework in several directions. First, integrating into the model pipeline dimensionality reduction techniques, such as principal component analysis or autoencoders, could further optimize the system’s performance in high-dimensional spaces, particularly in scenarios where the feature set is large or sparsely populated. Additionally, exploring alternative feature extraction methods, such as graph neural networks, may enhance the model’s ability to capture complex relationships within data, especially for datasets with inherent hierarchical or graph-based structure. Another promising avenue is extending ClustRecNet to optimize the parameter settings for each clustering algorithm recommended, possibly using reinforcement learning, thus widening its applicability and enhancing precision.
Declaration of generative AI and AI-assisted technologies in the manuscript preparation process
During the preparation of this work, the authors used Gemini in order to improve the language and readability of the manuscript. After using this tool/service, the authors reviewed and edited the content as needed and take full responsibility for the content of the published article.
CRediT authorship contribution statement
Mohammadreza Bakhtyari: Writing – original draft, Methodology, Software, Visualization, Investigation, Conceptualization, Data curation. Bogdan Mazoure: Writing – review and editing, Validation, Methodology. Renato Cordeiro de Amorim: Writing – review and editing, Validation, Methodology. Guillaume Rabusseau: Writing – review and editing, Validation. Vladimir Makarenkov: Writing – review and editing, Conceptualization, Methodology, Validation, Supervision, Investigation.
Declaration of competing interest
We declare that there are no conflicts of interest to report.
Data availability
Code and data are available at:
https://github.com/mrbakhtyari/ClustRecNet
Acknowledgments
This work was supported by le Fonds Québécois de la Recherche sur la Nature et les Technologies [grant 371537] and the Natural Sciences and Engineering Research Council of Canada [grant 249644].